Jefferson Lab HALL A Logbook for July 99 - All EXCEPT Automatic Entries

  • Main INDEX

    Saturday

  • 23996 07/31/99 23:32 J. LeRose H2 target in
  • 23994 07/31/99 23:22 J. LeRose ep finished
  • 23992 07/31/99 23:18 x. jiang at run 620, spectrometer screen
  • 23988 07/31/99 22:34 J. LeRose ep starts
  • 23985 07/31/99 22:19 X. Jiang fpp chamber hv 7.1.4 tripped at the end of run 617.
  • 23982 07/31/99 21:44 x. jiang fpp hv setting in run 617
  • 23981 07/31/99 21:44 x. jiang run 617 spectrometer screen
  • 23979 07/31/99 21:39 X. Jiang run 616 fpp chamber 8.1.16 tripped.
  • 23976 07/31/99 21:30 J. LeRose 50 microamps
  • 23975 07/31/99 21:24 J. LeRose 30 microamps
  • 23972 07/31/99 21:05 J. LeRose Spot
  • 23965 07/31/99 20:46 J. LeRose News from MCC
  • 23964 07/31/99 20:28 J. LeRose Beam!
  • 23963 07/31/99 20:01 J. LeRose Pulsed Beam
  • 23962 07/31/99 19:21 J. LeRose Beam Update
  • 23961 07/31/99 17:37 J. LeRose Beam Energy measurement
  • 23960 07/31/99 17:33 J. LeRose Accelerator gun Problems
  • 23959 07/31/99 16:54 R. Michaels small problems noted
  • 23958 07/31/99 16:37 J. LeRose Swing Shift Starts
  • 23957 07/31/99 16:16 X. Jiang Check FPP chamber HV .
  • 23955 07/31/99 15:59 kuss spectrometers ready
  • 23954 07/31/99 15:58 kuss no beam
  • 23953 07/31/99 14:50 kuss CHL recovery
  • 23952 07/31/99 11:23 kuss CHL, beam off
  • 23951 07/31/99 11:22 kuss magnets polarity change
  • 23945 07/31/99 09:30 kuss unrastered beam
  • 23941 07/31/99 08:55 kuss angle marks
  • 23940 07/31/99 08:34 kuss run during EP
  • Friday

  • 23915 07/30/99 16:50 R. Michaels info on trigger (read it)
  • 23913 07/30/99 16:15 M Jones second thoughts about epicsLogger
  • 23908 07/30/99 16:02 M Jones modified epicsLogger
  • 23907 07/30/99 15:50 M Jones fixed problem with runend.tcl
  • 23895 07/30/99 12:49 R. Michaels trigger status
  • 23890 07/30/99 11:52 S Strauch,R E-arm aerogel
  • 23885 07/30/99 06:56 X. Jiang raster current display on the scope not working
  • 23883 07/30/99 06:44 X Jiang run 579. c12 target
  • 23881 07/30/99 06:34 X Jiang Q3 quenched.
  • 23880 07/30/99 06:19 A. Deur eP measurement
  • 23878 07/30/99 05:07 X. Jiang run 578, c12 target.
  • 23876 07/30/99 05:03 X. Jiang run 577. Junk.
  • 23872 07/30/99 01:19 X. Jiang Spectrometer screen.
  • Thursday

  • 23868 07/29/99 22:05 R. Michaels re: E-arm CAMAC crate
  • 23862 07/29/99 20:29 R. Michaels CAMAC power supply
  • 23861 07/29/99 20:21 R. Michaels MLU problems on E-arm
  • 23860 07/29/99 10:22 mitchell Hall C BCM Signal and TOF RF
  • Wednesday

  • 23855 07/28/99 00:57 X Jiang H-arm Q3 quenched
  • 23854 07/28/99 00:14 X Jiang start shift at 00:00, beam is 9.5 uA,
  • Tuesday

  • 23849 07/27/99 17:28 Sabatie Beam back for Compton tests
  • 23844 07/27/99 10:58 tieulent linearity test
  • 23835 07/27/99 09:56 tieulent spot++
  • 23831 07/27/99 09:44 Tieulent Magnets ready
  • 23830 07/27/99 09:39 Wijesooriya end of sieve slit runs
  • 23829 07/27/99 09:21 Tieulent Set back magnet to 0%
  • 23826 07/27/99 09:01 Tieulent Magnet ready
  • 23825 07/27/99 08:53 averett HV change
  • 23824 07/27/99 08:42 Tieulent Change magnet field
  • 23820 07/27/99 08:19 Tieulent Electron set to +3.0%
  • 23819 07/27/99 08:10 Dale change electron arm field
  • 23818 07/27/99 08:08 Dale Good run
  • 23815 07/27/99 07:45 Dale ROC14 reboot
  • 23814 07/27/99 07:36 Dale beam back
  • 23811 07/27/99 06:01 Dale lost beam
  • 23809 07/27/99 05:52 Dale run 559
  • 23807 07/27/99 05:37 X. Jiang run 557, sieve slit pattern
  • 23804 07/27/99 05:30 Dale Run 558
  • 23803 07/27/99 05:25 Dale lost beam
  • 23800 07/27/99 05:20 Dale junk run
  • 23797 07/27/99 05:09 Dale magnet screen
  • 23795 07/27/99 05:05 dale spot looks good
  • 23793 07/27/99 05:02 dale checking spot
  • 23792 07/27/99 04:56 dale checking spot
  • 23789 07/27/99 04:36 dale finished Moller
  • 23788 07/27/99 04:06 Eugene Moller tests
  • 23787 07/27/99 00:54 dale scintillator trigger
  • 23786 07/27/99 00:39 dale sieve slits in
  • 23785 07/27/99 00:25 dale changing spectrometer fields
  • Monday

  • 23784 07/26/99 23:36 dale Moller tests
  • 23783 07/26/99 23:27 R. Michaels long Q^2 run 553
  • 23780 07/26/99 23:19 R. Michaels run 552 -- Q^2 measurement
  • 23777 07/26/99 23:11 R. Michaels alignment check
  • 23770 07/26/99 22:46 R. Michaels lead target runs
  • 23750 07/26/99 20:42 R. Michaels spot adjustment
  • 23744 07/26/99 20:27 R. Michaels Pb target tests -- continue
  • 23741 07/26/99 20:10 Templon beam is back
  • 23740 07/26/99 19:59 Templon no beam, reasons, BCM trips
  • 23738 07/26/99 15:44 R. Michaels sieve, Q^2 summary of weekend
  • 23737 07/26/99 14:31 D. Armstrong beam current limitations
  • 23736 07/26/99 14:24 R. Michaels beam back - spot check
  • 23733 07/26/99 11:37 R. Michaels no elastic lead seen
  • 23732 07/26/99 11:01 R. Michaels H-arm voltages turned off
  • 23731 07/26/99 10:49 R. Michaels Run 534 -- 8 micro on lead
  • 23727 07/26/99 10:42 R. Michaels run 533 -- 6 microamps
  • 23724 07/26/99 10:34 R. Michaels 3 microamps - run 532
  • 23719 07/26/99 10:16 R. Michaels temperature at 1 microamp
  • 23718 07/26/99 10:13 R. Michaels rates at 1 microamp
  • 23716 07/26/99 10:06 R. Michaels moving in lead target
  • 23715 07/26/99 09:52 R. Michaels spot for lead tests
  • 23706 07/26/99 09:30 R. Michaels setup for lead tests
  • 23704 07/26/99 07:52 Minsuk spot size
  • 23701 07/26/99 07:05 Minsuk spot size
  • 23697 07/26/99 03:56 dale raster change
  • 23698 07/26/99 03:56 Minsuk spot size
  • 23696 07/26/99 03:55 dale changed raster
  • 23691 07/26/99 02:37 Minsuk spot size
  • 23690 07/26/99 02:19 dale carbon elastic
  • 23687 07/26/99 02:16 dale analyzed run 520
  • 23685 07/26/99 02:02 dale increase current
  • 23684 07/26/99 01:54 Minsuk spot size
  • 23682 07/26/99 01:53 dale short run carbon elastic
  • 23680 07/26/99 01:46 Minsuk spot size
  • 23679 07/26/99 01:44 dale spot check
  • 23676 07/26/99 01:43 dale beam spot check
  • 23674 07/26/99 01:40 dale carbon target
  • 23673 07/26/99 01:36 dale collimators put in
  • 23671 07/26/99 01:05 dale magnet screen, run 517
  • 23669 07/26/99 01:00 dale optics studies, delta scan
  • 23667 07/26/99 00:45 dale unstable beam current
  • 23666 07/26/99 00:34 dale optics studies, delta scan
  • 23664 07/26/99 00:23 dale changed electron arm momentum
  • 23661 07/26/99 00:15 dale optics studies, hydrogen
  • Sunday

  • 23655 07/25/99 23:38 X. jiang magnet screen at the end of run 513
  • 23654 07/25/99 23:30 X. jiang run 513 H-arm only run.
  • 23651 07/25/99 23:03 X. jiang run 512, E-arm only run.
  • 23648 07/25/99 22:59 X. jiang magnet screen at the end of run 511
  • 23647 07/25/99 22:50 X. jiang run 511. H-arm 3.00386 Gev
  • 23645 07/25/99 22:46 X. jiang magnet screen at the end of run 510
  • 23644 07/25/99 22:36 X. jiang run 510, e-arm only,
  • 23641 07/25/99 22:18 X. jiang run 509 setting.
  • 23639 07/25/99 21:31 R. Michaels ...but NOT hadron arm !
  • 23638 07/25/99 21:25 R. Michaels suggest C12 elastic
  • 23634 07/25/99 20:12 X. jiang beam back at 20:10
  • 23633 07/25/99 20:02 X. Jiang e-arm Cherenkov turned on
  • 23632 07/25/99 18:14 X. jiang MCC called for system check.
  • 23630 07/25/99 18:00 X. jiang 17:55 put in 15cm H target,
  • 23627 07/25/99 17:47 X. jiang Asked mcc to set raster size to 2.0 (x) by 2.0 (y
  • 23626 07/25/99 17:45 R. Michaels recommended max H-arm field
  • 23621 07/25/99 17:35 X. jiang raster check
  • 23618 07/25/99 17:26 X. jiang run 503, the same as in run 501
  • 23615 07/25/99 17:17 X. jiang VDC drift time spectra looks not too good.
  • 23614 07/25/99 17:13 X. jiang run 501, sieve slit pattern as seen from H-arm.
  • 23613 07/25/99 17:11 K. Aniol un 502 setting the same as run 501.
  • 23611 07/25/99 17:06 K. Aniol e-arm mag. field values
  • 23610 07/25/99 17:05 X. jiang run 501 sieve slit holes .
  • 23608 07/25/99 16:33 X. jiang 15 uA on C12
  • 23604 07/25/99 15:43 D. Prout More Delta Scan/Optics summary
  • 23601 07/25/99 15:22 D. Prout 15 micro amp
  • 23595 07/25/99 13:13 Adaq -3% run carbon sieve done ?
  • 23593 07/25/99 12:27 Adaq carbon, sieve now in
  • 23590 07/25/99 11:25 Adaq some beam
  • 23589 07/25/99 11:23 Adaq Optics summary
  • 23588 07/25/99 11:10 Adaq Hall A still waiting
  • 23587 07/25/99 11:09 Adaq magnet change
  • 23586 07/25/99 11:03 Adaq run plan update
  • 23585 07/25/99 10:45 Adaq beam to hall a
  • 23582 07/25/99 09:30 Adaq bcm trips; hall A in the dog house
  • 23581 07/25/99 09:01 Adaq MCC test
  • 23580 07/25/99 08:00 Templon shift summary
  • 23579 07/25/99 07:56 Dale bad beam
  • 23576 07/25/99 07:14 Dale increased beam current
  • 23575 07/25/99 07:04 Dale delta scan, optics studies
  • 23571 07/25/99 06:28 Templon cryotarget alarms
  • 23570 07/25/99 05:52 Dale seive slit, delta scan, optics studies,carbon
  • 23568 07/25/99 05:40 Dale seive slit
  • 23567 07/25/99 05:36 Dale question, wire chambers
  • 23565 07/25/99 05:26 Templon cryotarget, loop 1
  • 23563 07/25/99 04:48 Dale delta scan, optics studies
  • 23561 07/25/99 04:21 Dale delta scan, optics studies
  • 23558 07/25/99 03:57 Dale optics studies spot size
  • 23555 07/25/99 03:43 Dale raster, spot adjust
  • 23553 07/25/99 03:21 Templon tgt change, trig change
  • 23551 07/25/99 02:52 Dale background measurement
  • 23548 07/25/99 02:22 Dale Q^2 measurement
  • 23543 07/25/99 00:11 R. Michaels alignment check
  • Saturday

  • 23526 07/24/99 17:59 R. Michaels re: optics run plan
  • 23524 07/24/99 14:29 D. Margaziot No polarized beam
  • 23523 07/24/99 12:26 R. Michaels plan for optics calibration, Pb test
  • 23522 07/24/99 11:55 R. Michaels run plan -- thermionic
  • Friday

  • 23510 07/23/99 15:29 Humensky update from MCC
  • 23509 07/23/99 15:11 Humensky cathode QE low ; beam stopper pressure fault
  • 23505 07/23/99 14:40 Humensky loop 2 helium pressure went low; okay now
  • 23504 07/23/99 14:35 Humensky TACO change
  • 23503 07/23/99 14:34 Humensky beam returns -- 10uA -- TACO
  • 23499 07/23/99 13:30 markowitz spot++
  • 23492 07/23/99 12:11 bogdanw Moller Q trip
  • 23491 07/23/99 11:47 Humensky changed data disk
  • 23488 07/23/99 11:04 rsholmes EPICS
  • 23487 07/23/99 10:33 Humensky Q3H trip ; He briefly low
  • 23486 07/23/99 10:21 Humensky temperature rising at top of loop1
  • 0 07/23/99 10:20 no_name no_key
  • 23483 07/23/99 09:44 Humensky beam modulation off for run 5508
  • 23482 07/23/99 09:42 Humensky beam quiet; minirun length back to 2500pairs
  • 23481 07/23/99 09:36 Humensky TACO change
  • 23480 07/23/99 09:25 R. Michaels meaningless EPICS errors on happex1
  • 23479 07/23/99 09:22 markowitz spot++ screen shot
  • 23474 07/23/99 09:06 Humensky TACO change
  • 23470 07/23/99 08:02 Templon cryotarget and ESR puzzles
  • 23468 07/23/99 07:26 Templon TACO change
  • 23463 07/23/99 06:30 Templon TACO voltage change
  • 23462 07/23/99 06:22 Templon TACO change
  • 23448 07/23/99 01:43 G. Rutledge Taco
  • 23445 07/23/99 00:24 G. Rutledge Injector DAQ Hangup
  • Thursday

  • 23440 07/22/99 23:33 markowitz rf trip
  • 23436 07/22/99 22:37 humensky TACO changes
  • 23433 07/22/99 21:53 humensky late entry--PITA slope measurements
  • 23429 07/22/99 21:32 markowitz first production data of shift
  • 23402 07/22/99 19:03 markowitz spot screen shot
  • 23399 07/22/99 18:56 markowitz pita slope
  • 23395 07/22/99 18:34 humensky PITA slope still bad
  • 23358 07/22/99 13:26 averett beam off
  • 23357 07/22/99 13:10 Humensky controlled access to realign Pockels Cell
  • 23356 07/22/99 12:36 averett spot++
  • 23353 07/22/99 11:41 Eugene Moller results
  • 23352 07/22/99 10:24 averett Wein spin angle
  • 23347 07/22/99 06:27 G. Rutledge No PITA
  • 23339 07/22/99 05:34 G. Rutledge Still no PITA
  • 23331 07/22/99 04:01 Humensky high beam intensity noise
  • 23307 07/22/99 01:18 Humensky happex1 running slow; rebooted
  • 23304 07/22/99 00:21 G. Rutledge Raster Shape... o.k.
  • Wednesday

  • 23293 07/21/99 23:46 Markowitz spot
  • 23288 07/21/99 20:27 Markowitz spot++ screen shot
  • 23283 07/21/99 08:49 X. Jiang FPP HV monitor summary screen.
  • 23282 07/21/99 08:48 X. Jiang FPP HV screen, card-8 ch 9- 13. and ch 16.
  • 23281 07/21/99 08:46 X. Jiang FPP HV screen. Card-8.
  • 23279 07/21/99 08:44 X. Jiang FPP HV screen, card 7
  • 23280 07/21/99 08:44 X. Jiang FPP HV card 7, ch 8-13.
  • Tuesday

  • 23276 07/20/99 14:46 rsholmes Baby correction
  • 23272 07/20/99 14:27 X. Jiang Hall-A HV control.
  • 23271 07/20/99 14:25 X. Jiang FPP low voltage
  • 23270 07/20/99 14:22 X. Jiang Magnet screen
  • 23265 07/20/99 11:23 Jiang FPP chamber on
  • 23263 07/20/99 03:13 G. Rutledge Background Scans Complete
  • Monday

  • 23238 07/19/99 23:30 R. Michaels scan - last of spectr. DAQ
  • 23233 07/19/99 22:45 R. Michaels field scan runs
  • 23228 07/19/99 22:07 R. Michaels thresholds, H-arm VDC, scan
  • 23225 07/19/99 21:58 Humensky beam noise -- late entry
  • 23224 07/19/99 21:52 R. Michaels more dipole scan runs
  • 23218 07/19/99 21:32 R. Michaels runs for field scan
  • 23213 07/19/99 21:17 R. Michaels dipole scan about to start
  • 23211 07/19/99 21:00 R. Michaels run 407 -- beam right 2 mm
  • 23208 07/19/99 20:49 R. Michaels beam centered again
  • 23205 07/19/99 20:39 R. Michaels run 405, position scan
  • 23202 07/19/99 20:30 R. Michaels run 404 -- spot left
  • 23200 07/19/99 20:25 R. Michaels results of empty target
  • 23199 07/19/99 20:15 R. Michaels dummy target, sieve slit run 403
  • 23197 07/19/99 20:08 R. Michaels dummy target run -- step 4
  • 23189 07/19/99 19:49 R. Michaels run 398 -- step 2
  • 23184 07/19/99 19:38 R. Michaels run 397 to check alignment
  • 23181 07/19/99 19:34 R. Michaels raster adjustment
  • 23176 07/19/99 19:18 R. Michaels Q^2, backgrounds - run plan
  • 23165 07/19/99 16:31 Adev good beam tune
  • 23162 07/19/99 16:18 Adev spot++ screen shot
  • 23130 07/19/99 05:29 M.Baylac New tune : check coil pulsing
  • 23129 07/19/99 05:19 M.Baylac New beam tune
  • 23118 07/19/99 04:31 M.Baylac Beam back : check tune
  • 23087 07/19/99 02:36 Minsuk Spot check
  • 23081 07/19/99 02:01 Humensky access made to reset tripped moller quads
  • Sunday

  • 23079 07/18/99 22:43 Humensky figure for last entry
  • 23078 07/18/99 22:39 Humensky status update
  • 23050 07/18/99 08:42 Lhuillier Magnet recycling
  • 23032 07/18/99 06:45 M.Baylac Beam modulation and raster size
  • 23026 07/18/99 06:23 M.Baylac PITA slope
  • 23003 07/18/99 05:42 M.Baylac Test with beam modulation OFF
  • 22999 07/18/99 05:29 M.Baylac Raster size
  • Saturday

  • 22980 07/17/99 21:38 humensky Q3H quenched, fix imminent
  • 22979 07/17/99 20:18 humensky Update on CHL: est. 4 more hours
  • 22978 07/17/99 19:10 humensky updated detector pedestals
  • 22976 07/17/99 16:50 humensky new PITA slope and slug
  • 22975 07/17/99 16:49 humensky late entry: TACO corrections
  • 22974 07/17/99 15:21 M Jones Cold compressor down, CHL down
  • 22973 07/17/99 15:17 M Jones different elec dipole current for different probes
  • 22971 07/17/99 13:30 Lhuillier New tunning is OK!!
  • 22965 07/17/99 11:58 Calarco beam back (maybe)
  • 22962 07/17/99 09:08 Lhuillier Bad tunning- stop beam in HallA.
  • 22943 07/17/99 07:06 Minsuk Beam spot
  • 22920 07/17/99 06:22 Minsuk First beam spot after ~6 hours
  • 22899 07/17/99 00:14 R. Michaels Q^2 tests ended -- return to happex
  • Friday

  • 22893 07/16/99 23:40 R. Michaels prepare for empty target run
  • 22891 07/16/99 23:22 R. Michaels C12, sieve, scint trigger
  • 22888 07/16/99 23:13 R. Michaels spot at 3A, 3B during 299
  • 22887 07/16/99 22:59 R. Michaels Run 299 -- sieve with C12
  • 22885 07/16/99 22:50 R. Michaels wierd spot size - run 298
  • 22881 07/16/99 22:45 R. Michaels run 297 -- still wide in X
  • 22879 07/16/99 22:41 R. Michaels BABY IS BORN !!!
  • 22878 07/16/99 22:39 R. Michaels carbon target runs
  • 22873 07/16/99 22:23 R. Michaels run 294 results
  • 22871 07/16/99 22:10 R. Michaels run 294 - sieve slit run
  • 22870 07/16/99 22:07 R. Michaels spot for sieve slit tests
  • 22869 07/16/99 22:06 R. Michaels spot check with sieve slit
  • 22858 07/16/99 21:43 R. Michaels run 289 beam low
  • 22856 07/16/99 21:41 R. Michaels run 288 beam up Q^2
  • 22854 07/16/99 21:38 R. Michaels beam up 2 mm - run 288
  • 22852 07/16/99 21:32 R. Michaels spot to right -- Q^2
  • 22850 07/16/99 21:24 R. Michaels run 287 spot to right
  • 22848 07/16/99 21:21 R. Michaels run 286
  • 22845 07/16/99 21:13 R. Michaels run 285
  • 22842 07/16/99 20:50 R. Michaels beam problems in chicane
  • 22841 07/16/99 20:35 R. Michaels run 284
  • 22838 07/16/99 20:33 R. Michaels Q^2 from run 283
  • 22836 07/16/99 20:28 R. Michaels run 283, small spot
  • 22834 07/16/99 20:26 R. Michaels run 282 - beam centered
  • 22831 07/16/99 20:13 R. Michaels step 2 commencing
  • 22825 07/16/99 20:01 R. Michaels alignement ok -- run 279
  • 22824 07/16/99 19:58 R. Michaels run 279 -- repeat
  • 22822 07/16/99 19:52 R. Michaels spot size
  • 22819 07/16/99 19:36 R. Michaels run plan for tonight
  • 22805 07/16/99 15:28 D. Armstrong TACO adjust
  • 22799 07/16/99 14:46 D. Armstrong TACo changed
  • 22794 07/16/99 14:02 D. Armstrong more on coil pulsing
  • 22791 07/16/99 13:36 D. Armstrong coil pulsing turned back on
  • 22787 07/16/99 12:15 k. wijesoori target alarm not working
  • 22783 07/16/99 10:22 k. wijesoori injector problems
  • 22778 07/16/99 09:51 k. wijesoori Current modulation
  • 22777 07/16/99 09:31 K. Wijesoori Intencity modulation
  • 22771 07/16/99 09:11 k. wijesoori did not do ramping
  • 22770 07/16/99 09:07 B. Tonguc linearity check
  • 22751 07/16/99 04:58 Minsuk New TACO
  • 22736 07/16/99 03:46 Minsuk New TACO
  • 22733 07/16/99 03:02 Minsuk New TACO value
  • 22725 07/16/99 01:48 Cavata Position helicity differences
  • Thursday

  • 22702 07/15/99 22:56 J. Roche raster shape
  • 22692 07/15/99 22:15 Eugene Moller results
  • 22689 07/15/99 19:24 Humensky new PITA slope and other good news
  • 22673 07/15/99 15:43 Eugene Moller results
  • 22669 07/15/99 13:07 B. Tonguc Inj. DAQ crashed.
  • 22658 07/15/99 09:53 R. Michaels hac is back, sort-of
  • 22646 07/15/99 07:38 M.Baylac Current oscillations time constant
  • 22645 07/15/99 07:02 Finn Taco value runs 5395-5397
  • 22635 07/15/99 05:32 Finn Half wave plate angle readout
  • 22634 07/15/99 05:28 M.Baylac Beam is back!!!!
  • 22627 07/15/99 03:32 M.Baylac Remove target
  • 22626 07/15/99 02:59 M.Baylac Sign tests
  • Wednesday

  • 22615 07/14/99 22:25 Humensky MCC says they don't know how long they'll be down
  • 22613 07/14/99 14:29 mitchell Hall Off for Tour
  • 22599 07/14/99 09:36 mitchell Resume Runing at 8:32--Blind
  • 22598 07/14/99 09:04 A. Deur Final results of eP energy measurement of may 1999
  • 22592 07/14/99 08:13 R. Michaels problems with hac
  • 22590 07/14/99 05:39 M.Baylac Remove target
  • 22589 07/14/99 05:37 Finn Beam off due to HAC failure
  • 22588 07/14/99 05:29 M.Baylac HAC crashed
  • 22579 07/14/99 04:09 Finn taco for runs 5374-5382
  • 22553 07/14/99 00:27 M.Baylac Increase current
  • 22552 07/14/99 00:21 Humensky beam has gotten noisier
  • 22549 07/14/99 00:05 Humensky TACO correction
  • 22548 07/14/99 00:04 Humensky new PITA slope and slug
  • 22546 07/14/99 00:02 R. Michaels network problems
  • Tuesday

  • 22539 07/13/99 22:56 Kooijman new TACO
  • 22536 07/13/99 22:27 R. Michaels Inj run 3978 was a bust
  • 22535 07/13/99 22:16 R. Michaels NMR checked -- possible mag drift
  • 22534 07/13/99 22:06 Tonguc new TACO value
  • 22529 07/13/99 21:39 Eugene A script to monitor the HAPPEX detectors HV
  • 22528 07/13/99 20:53 Tonguc new TACO
  • 22526 07/13/99 20:35 R. Michaels spare also dropped in run 5366
  • 22524 07/13/99 20:09 R. Michaels detector asy for run 5367
  • 22523 07/13/99 19:49 Eugene Moller results , spin dance results
  • 22522 07/13/99 19:45 R. Michaels mysterious drops in det2
  • 22519 07/13/99 19:08 Tonguc new TACO
  • 22516 07/13/99 17:50 R. Michaels beam is back
  • 22508 07/13/99 16:48 mitchell Shift Summary
  • 22506 07/13/99 15:44 rsholmes Software change
  • 22505 07/13/99 13:19 mitchell Adjust JT loop 3
  • 22504 07/13/99 13:13 mitchell Moller "mini dance" starts
  • 22498 07/13/99 12:15 rsholmes EPICS changes
  • 22496 07/13/99 12:04 mitchell Fast Feedback Studies
  • 22489 07/13/99 11:10 mitchell Injector Daq dies
  • 22483 07/13/99 10:09 mitchell Hall C down for day
  • 22471 07/13/99 07:08 Finn Taco value run 5355
  • 22468 07/13/99 06:19 Finn Taco values runs 5353-5354
  • 22458 07/13/99 04:01 Finn taco values runs 5349-5352
  • 22442 07/13/99 01:08 Humensky Hall C bcm5 back
  • 22438 07/13/99 00:38 Humensky Hall C charge asymmetry
  • 22436 07/13/99 00:34 M.Baylac Reboot ROC 14
  • 22432 07/13/99 00:11 Humensky sweet spot done; back in production
  • Monday

  • 22419 07/12/99 20:49 Eugene Moller results
  • 0 07/12/99 19:37 no_name no_key
  • 22404 07/12/99 14:32 mitchell pulsed beam soon
  • 22403 07/12/99 14:30 mitchell Target Parked at Low F
  • 22401 07/12/99 13:28 mitchell Beam in 1 hr, Again
  • 22400 07/12/99 13:27 mitchell Empty
  • 22399 07/12/99 12:32 mitchell Pockels Cell HV switch
  • 22398 07/12/99 10:17 rsholmes injector DAQ not working
  • 22397 07/12/99 08:53 R. Michaels happex1 MSS backups ok now
  • 22395 07/12/99 07:42 M.Baylac Time estimate before beam comes back
  • 22394 07/12/99 05:02 M.Baylac No beam at all
  • 22393 07/12/99 01:12 M.Baylac P-gun pbs update
  • 22392 07/12/99 01:01 M.Baylac P-Gun pb
  • 22388 07/12/99 00:41 Younus New TACO value
  • Sunday

  • 22373 07/11/99 22:44 Tonguc new TACO value
  • 22369 07/11/99 21:31 Tonguc new TACO value
  • 22366 07/11/99 20:30 Tonguc new TACO
  • 22363 07/11/99 19:21 Tonguc new TACO
  • 22360 07/11/99 18:43 Humensky changed feedback minirun length and asymmetry cutpoint
  • 22359 07/11/99 18:18 Humensky new PITA slope
  • 22358 07/11/99 18:17 B. Tonguc new PMT voltage
  • 22353 07/11/99 17:57 B. Tonguc Four Harp Scan
  • 22346 07/11/99 15:19 R. Michaels controlled access
  • 22345 07/11/99 14:58 R. Michaels beam off -- injector work
  • 22341 07/11/99 14:23 R. Michaels beam back, run 5330
  • 22337 07/11/99 14:19 R. Michaels spare det HV on H-arm had been DISABLED
  • 22336 07/11/99 14:01 R. Michaels beam off for injector tuning
  • 22332 07/11/99 13:08 R. Michaels beam noise problem
  • 22331 07/11/99 12:43 R. Michaels beam back, but unstable (ped calib ?)
  • 22329 07/11/99 12:03 R. Michaels injector DAQ crash / pulsed beam
  • 22325 07/11/99 11:42 R. Michaels large beam noise noticed
  • 22322 07/11/99 11:02 R. Michaels injector DAQ ok now
  • 22321 07/11/99 10:40 R. Michaels injector DAQ not working
  • 22317 07/11/99 09:24 R. Michaels TACO change
  • 22311 07/11/99 07:56 Humensky TACO corrections owl shift
  • 22299 07/11/99 04:29 Humensky hall C to 60uA
  • 22296 07/11/99 03:43 Humensky note about online BCM5 asymmetry reports
  • 22295 07/11/99 03:24 Humensky New TACO slope
  • 22294 07/11/99 03:16 Humensky Q3 hadron is back
  • 22288 07/11/99 01:48 Humensky Q3 hadron quenched
  • 22287 07/11/99 01:34 Humensky Return of Hall C begins
  • Saturday

  • 22243 07/10/99 15:59 R. Michaels magnets epics screen
  • 22240 07/10/99 14:41 R. Michaels happex1 cannot back up to MSS
  • 22237 07/10/99 13:51 R. Michaels spot size, production mode
  • 22236 07/10/99 13:36 rsholmes Injector modulation
  • 22229 07/10/99 13:21 R. Michaels oversampling tests
  • 22223 07/10/99 13:09 R. Michaels boiling tests
  • 22198 07/10/99 11:58 R. Michaels upstream coil pulsing, now Compton
  • 22189 07/10/99 11:10 Eugene Moller results
  • 22188 07/10/99 10:12 R. Michaels wrong Wein setting !
  • 22186 07/10/99 08:24 R. Michaels access to reset Moller quad
  • 22185 07/10/99 08:18 R. Michaels status; Moller starting
  • 22176 07/10/99 05:48 finn ep energy measurement
  • 22164 07/10/99 02:29 Humensky starting production running (hopefully)
  • 22160 07/10/99 01:19 finn spot shape is now 3.5 by 4.5
  • Friday

  • 22142 07/09/99 23:45 Tieulent Beam is back
  • 22090 07/09/99 10:21 R. Michaels Q^2 Measurement (and T0)
  • 22088 07/09/99 09:52 R. Michaels attenuation checkout
  • 22087 07/09/99 09:44 R. Michaels alignment check
  • 22084 07/09/99 09:30 R. Michaels alignment check num 1
  • 22082 07/09/99 09:20 R. Michaels Detector checks
  • 22081 07/09/99 09:14 R. Michaels run 137 -- elastic
  • 22079 07/09/99 09:06 R. Michaels HV on. New B fields
  • 22077 07/09/99 08:45 R. Michaels spot size 2 mm square
  • 22070 07/09/99 08:23 R. Michaels plan for alignment, etc
  • 22065 07/09/99 07:43 roblin ep measurement
  • 22059 07/09/99 07:01 jiang Coda run 130,junk
  • 22054 07/09/99 06:07 Saha Arc Energy Measurement
  • Thursday

  • 22037 07/08/99 22:53 Saha, Jiang OTR study
  • 22030 07/08/99 21:02 Tieulent Setup Energy Measurment
  • 22029 07/08/99 20:43 K. McCormick Target GUI changes
  • Wednesday

  • 22009 07/07/99 18:26 K. McCormick High Voltage GUIs updated
  • Tuesday

    
    

    Monday

    
    

    Friday